# references.bib

### v3

@@ Abstract
In this paper we provide an explicit construction of the left adjoint to the inclusion of Heyting algebras in the category of distributive lattices using duality, generalizing a technique due to Ghilardi.This is employed to give a new concrete description of colimits and free Heyting algebras. Using such tools, we prove some results about the category of Heyting algebra. These are both known, concerning the amalgamation property and related facts; and new, such as the fact that the category of Heyting algebras is co-distributive. We also study some generalizations and variations of this construction to different related settings. We analyse some subvarieties of Heyting algebras – such as Boolean algebras, $\mathsf{KC}$ and $\mathsf{LC}$ algebras, and show how the construction can be adapted to such cases. Second, we study the relationship between the category of image-finite posets with p-morphisms and the category of posets with monotone maps, and using a variation of the above ideas, provide an adjunction between such categories.
@@ 1 Introduction
In algebraic settings, many problems boil down to the analysis of free objects: free groups, free modules over a ring, polynomial rings, amongst others. This ubiquity stems from the fact that understanding free objects of this kind allows a very fine combinatorial and categorical understanding of the structure of the classes of algebras of interest. As a result, they can encode a lot of the axiomatic structure of such algebras, as well as provide access to key theorems concerning finitely presented, finitely generated, or projective algebras, all of which intimately depend on free ones. Such algebras are always possible to describe using “word constructions”, or equivalently, by computing colimits of the algebras in terms of generators and relations, but such representations are often too abstract to yield useful results. As such, making them more concrete presents an attractive and often important problem.
In some algebraic settings, this concreteness can be obtained by duality theory. In its origins, in Stone’s celebrated theorem [Stone1936], this ‘duality’ associates to each given Boolean algebra $\mathcal{B}$ a unique compact, Hausdorff and zero-dimensional space $X_{B}$, corresponding to the prime spectrum of the Boolean algebra. The existence of such a duality makes several constructions much simpler: for instance, to calculate coproducts of Boolean algebras, one need only describe products of Stone spaces, which turn out to be exactly the usual topological products. Since arbitary free algebras can be computed as coproducts of $1$-generated algebras, this also allows an easy computation of free Boolean algebras.
A similar and more general situation holds concerning distributive lattices and their prime spectra. Here, with a given distributive lattice $\mathcal{D}$, one takes the prime spectrum $X_{D}$ equipped with an order $\leq_{D}$, such that the pair $(X_{D},\leq_{D})$ forms a so-called Priestley space [Priestley1970]. Here again the calculation of coproducts simplifies, dually, to the calculation of products of Priestley spaces, which are simple set-theoretic constructions.
The case of interest in this paper is that of Heyting algebras. These are objects in the signature $(H,\wedge,\vee,\rightarrow,0,1)$, where $(H,\wedge,\vee,0,1)$ is a bounded lattice, and where the meet operation $\wedge$ has an adjoint, i.e., for $a,b,c\in H$,
$a\wedge b\leq c\iff a\leq b\rightarrow c.$
These structures have been studied extensively in several different mathematical contexts, where they arise naturally: 1) they are algebraic models for intuitionistic logic [Chagrov1997-cr]; 2) all lattices of open sets of topological spaces are Heyting algebras [McKinsey1944]; 3) locales are complete Heyting algebras [PP2012]; 4) In categorical logic they arise as the decategorification of cartesian closed categories; 5) In topos theory they arise as the lattices of subobjects of toposes; amongst others. Following the work of Esakia [Esakiach2019HeyAlg], one can also represent the spectrum of a Heyting algebra in a way that accounts for the implication connective using so-called “Esakia spaces”.
Despite having such a good representation, the structure of free Heyting algebras is much more difficult than that of Boolean algebras or distributive lattices: already the free algebra on one generator is infinite, observed by Rieger [Rieger1952-RIEOTL] and Nishimura [nishimuralattice]. The free algebra on two generators is notoriously difficult, having no easy lattice theoretic description.
@@
    The construction of the $n$-universal model (due independently to Bellissima [Bellissima1986], Grigolia [Grigolia1995]) and Shehtman [shehtmanrieger], which clarified the size and structure of irreducible elements of the free $n$-generated algebra, as well as some of its properties like completeness;
@@
Despite the importance of such advances, the situation presently in the literature presents a big contrast between what is known about Boolean algebras, distributive lattices and Heyting algebras: the former have descriptions of their free algebras, coproducts, and free one-step extensions for any number of generators, and any original algebra, whilst the latter are only available for very limited situations, involving finite algebras.
The main contribution of this paper is to provide a generalization of the constructions presented by Ghilardi to the infinite case. The main hurdle in this is realizing wherein finiteness plays a role, to allow using Priestley duality in place of Birkhoff duality for finite distributive lattices, a task which turns out to not be obvious. This turns out to amount to describing, in dual terms, the left adjoint to the inclusion of $\mathbf{HA}$, the category of Heyting algebras, in $\mathbf{DL}$, the category of distributive lattices.
Having such a description at hand, we employ such a construction for several purposes:
@@
    We provide a description of the free Heyting algebras on any number of generators;
@@
    We provide an associated concrete description of the coproducts of Heyting algebras, as well as that of the pushout of Heyting algebras. The latter is in turn used to give a new proof to the well-known fact that the category of Heyting algebras has the amalgamation property.
@@
    Using the construction of coproducts we prove a new result showing that the category of Heyting algebras is co-distributive, as well as several related properties concerning co-extensivity of this category.
We then proceed to extend these results in several directions, staying in the realm of Heyting algebras and related categories. We first show that this construction can be adapted when dealing with some subvarieties of Heyting algebras. Rather than providing a general theory, we focus on three important and illustrative cases: Boolean algebras, $\mathsf{KC}$ algebras and $\mathsf{LC}$ algebras. Finally, we look at the relationship between the category of posets with p-morphisms and the category of posets with monotone maps, and provide a right adjoint to the inclusion which is heavily inspired by the above constructions. The latter has been used in joint work with Nick Bezhanishvili [coalgebraicintuitionisticmodal] to address an open problem concerning coalgebraic representations of intuitionistic modal algebras. We also allude to further results on the category of Heyting algebras which we plan to develop in a sequel to this paper [openmapsandinterpolantcomplexity].
The structure of the paper is as follows: in Section 2 we recall some basic preliminaries, and fix some notation moving forward. Section 3 contains the main technical tools, in the form of the description of the free Heyting algebra generated by a distributive lattice preserving a given set of relative pseudocomplements; in Section 4 we showcase some basic applications of the aforementioned theory to the study of Heyting algebras: some well-known results, like the amalgamation theorem, are shown here through direct methods, as well as some seemingly new results, like the co-distributivity of the category of Heyting algebras. In Section 5 we look at the case of subvarieties of Heyting algebras and illustrate how these constructions can be fruitfully analysed there. In Section 6 we look at the corresponding relationship between the category of posets with p-morphisms and the category of posets with monotone maps. We conclude in Section 7 with some outlines of future work.
@@ 2 Preliminaries
Proposition 7.
If $X$ is an arbitrary set, then $\mathbf{D}(X)$ is isomorphic to the free distributive lattice on $X$ many generators. Moreover, if $X$ is finite, then $F_{D}(X)$ is isomorphic to the discrete topology on on $(\mathcal{P}(X),\supseteq)$.
@@
We also recall the well-known construction of the Vietoris hyperspace, due to Leopold Vietoris [Vietoris1921] (see also [guramremarksonhyperspacesforpriestleyspaces]): let $(X,\leq)$ be a Priestley space. Let $V(X)$ be the set of all closed subsets of $X$. We give this set a topology by considering a subbasis consisting, for $U,V\subseteq X$ clopen sets:
@@
This is sometimes called the “hit-and-miss topology”, and the resulting space is called the Vietoris hyperspace. On this space we define an order relation $C\preceq D$ if and ony if [^1] The reason this order is reverse inclusion has to with the choice to use upsets rather than downsets; in [ghilardifreeheyting], the author uses downsets, which is why the swap no reversion is needed. $D\subseteq C$. Then we have the following:
Lemma 9.
The space $(V(X),\preceq)$ is a Priestley space.
@@
Otherwise, we have that $C\neq\emptyset$, so $C\in V(X)$, and since $C\notin\langle V_{j}\rangle$ for each $i$, by definition $C\in[U_{i}]$ for some $i$. Since $X-\bigcup_{i\in I}V_{i}\subseteq U_{i}$ then $X-U_{i}\subseteq\bigcup_{i\in I}V_{i}$, so by compactness, $X-U_{i}\subseteq V_{i_{0}}\cup...\cup V_{i_{n}}$. Hence consider the finite subcover $\{[U_{i}]\}\cup\{\langle V_{i_{j}}\rangle:j\in\{1,...,n\}\}$. If $D$ is any closed set, and $D\subseteq U_{i}$, then we are done; otherwise, $D\cap X-U_{i}\neq\emptyset$, so $D\cap V_{i_{0}}\cup...\cup V_{i_{n}}\neq\emptyset$, from whence the result follows.
Now we prove the Priestley Separation axiom. Assume that $C\npreceq D$ are closed subsets. Since the space $X$ is a Stone space, we know that $C=\bigcap_{i\in I}V_{i}$ and $D=\bigcap_{j\in J}W_{j}$ where $V_{i},W_{j}$ are clopen. Hence there must be some $i$ such that $C\subseteq V_{i}$ and $D\nsubseteq V_{i}$; i.e., $C\in[V_{i}]$ and $D\notin[V_{i}]$. Since $[V_{i}]$ is clearly a clopen upset under this order, we verify the Priestley separation axiom. ∎
@@ 3.1 Conceptual Idea of the Construction and the Finite Case
    $(a\vee b)\rightarrow c=a\rightarrow c\wedge b\rightarrow c$;
@@
    $a\rightarrow(b\wedge c)=a\rightarrow b\wedge a\rightarrow c$;
@@
    $a\rightarrow b\wedge b\rightarrow c\leq a\rightarrow c$.
@@ 3.2 G-Open Subsets and Vietoris Functors
Now assume that $p=f^{-1}$ is $g$-open, and $q=g^{-1}$. Assume that $f(x)\leq y$ where $x\in X$ and $y\in Y$. By duality, and abusing notation, this means that $p^{-1}[x]\subseteq y$. Consider $\mathsf{Fil}(x,\{p(q(a)):q(a)\in y\})$ and $\mathsf{Id}(\{p(q(b)):q(b)\notin y\})$. We claim these do not intersect. Because otherwise, for some $c\in x$, $c\wedge p(q(a))\leq p(q(b))$. Hence $c\leq p(q(a))\rightarrow p(q(b))$, since these exist and are preserved by $p$, and so $c\leq p(q(a)\rightarrow q(b))$. So $p(q(a)\rightarrow q(b))\in x$. Hence $q(a)\rightarrow q(b)\in p^{-1}[x]$, and so $q(a)\rightarrow q(b)\in y$, a contradiction. Hence by the Prime filter theorem (see Theorem 4), we can extend $x$ to a filter $x^{\prime}$ which does not intersect the presented ideal. By definition, working up to natural isomorphism, we then have that $g(f(x^{\prime}))=g(y)$, which was to show. ∎
@@
Throughout, fix $g:X\to Y$ a continuous and order-preserving map (the case for several such maps being preserved is entirely similar). Recall from Lemma 9 that $(V(X),\preceq)$ is a Priestley space. From this space we can move closer to the space we will be interested, by first considering $V_{r}(X)\subseteq V(X)$, the space of rooted closed subsets, with the induced order and the subspace topology. On this we can prove the following [^2] This is stated without proof in [Bezhanishvili2011heytingcoalgebra, Lemma 6.1]. The key idea of the proof below was communicated to me by Mamuka Jibladze.:
Lemma 13.
@@
Finally, we will refine the rooted subsets to the ones we are truly interested in – the $g$-open ones. The following lemma encapsulates the key fact about this.
Lemma 14.
@@
Assume that $M$ is not $g$-open. This means that if $x$ is the root of $M$, there is some $y$ such that $x\leq y$, but for each $k\in M$, $g(k)\neq g(y)$. Hence $g(y)\notin g[M]$, and hence $\{g(y)\}\cap g[M]=\emptyset$. Since $g$ is a continuous map between Stone spaces, $g[M]$ is closed; hence, since this is a Priestley space, there is a clopen set of the form $-U_{i}\cup V_{i}$ for $U_{i},V_{i}$ clopen upsets, such that $g[M]\subseteq-U_{i}\cup V_{i}$, and $g(y)\notin-U_{i}\cup V_{i}$. Then we have that $M\subseteq-g^{-1}[U_{i}]\cup g^{-1}[V_{i}]$ and $M\cap{\downarrow}(g^{-1}[U_{i}]-g^{-1}[V_{i}])\neq\emptyset$. Hence consider
@@
Because $X$ has relative pseudocomplements of sets of the form $g^{-1}[U]$, by Proposition 5 we have that ${\downarrow}(g^{-1}[U_{i}]-g^{-1}[V_{i}])$ will be clopen in $X$. Hence the set $S$ is an intersection of subbasic sets. Moreover, $M$ belongs there. Now if $N$ belongs there, then first $N\subseteq-g^{-1}[U_{i}]\cup g^{-1}[V_{i}]$; and also, by downwards closure, its root $k$ will be below some point $y^{\prime}$ in $g^{-1}[U_{i}]-g^{-1}[V_{i}]$. But then for each $m\in N$, $g(m)\neq g(y^{\prime})$, so $N$ is not $g$-open.

### v4

@@ Abstract
In this paper we provide an explicit construction of the left adjoint to the inclusion of Heyting algebras in the category of distributive lattices using duality, generalizing a technique due to Ghilardi. This is employed to give a new concrete description of colimits and free Heyting algebras. Using such tools, we obtain in a uniform way some results about the category of Heyting algebra. These are both known, concerning the amalgamation property and related facts; and new, such as the fact that the category of Heyting algebras is comonadic over the category of distributive lattices. We also study some generalizations and variations of this construction to different related settings: (1) we analyse some subvarieties of Heyting algebras – such as Boolean algebras, $\mathsf{KC}$ and $\mathsf{LC}$ algebras, and show how the construction can be adapted to such cases, and deriving some logical properties of $\mathsf{LC}$ in the process; (2) we study an adjunction between Heyting algebras and Boolean algebras, showing how this provides a categorical semantics for inquisitive logic; (3) we study the relationship between the category of image-finite posets with p-morphisms and the category of posets with monotone maps, and using a variation of the above ideas, provide an adjunction between such categories, with applications in the coalgebraic approach to intuitionistic modal logic.
@@ 1 Introduction
Free algebras – in their form of Lindenbaum-Tarski algebras – play an important role in analyzing systems of non-classical logics [Chagrov1997-cr, Blackburn2002-fd, Galatos2007-vu]. For this purpose, duality theory has played a remarkable role, turning the abstract “word constructions” of free algebras into concrete and analyzable objects: for instance, to calculate coproducts of Boolean algebras, one need only describe products of Stone spaces, which turn out to be exactly the usual topological products. Since arbitary free algebras can be computed as coproducts of $1$-generated algebras, this also allows an easy computation of free Boolean algebras. A similar and more general situation holds concerning distributive lattices and their prime spectra.
In this paper we focus on the case of Heyting algebras, which are the algebraic models of $\mathsf{IPC}$, the intuitionistic propositional calculus, or intuitionistic logic. Despite having a good representation theory in terms of Esakia spaces, the structure of free Heyting algebras is much more difficult than that of Boolean algebras or distributive lattices: already the free algebra on one generator is infinite, observed by Rieger [Rieger1952-RIEOTL] and Nishimura [nishimuralattice]. The free algebra on two generators is notoriously difficult, having no, known, easy lattice theoretic description.
@@
    The construction of the $n$-universal model (due independently to Bellissima [Bellissima1986], Grigolia [Grigolia1995] and Shehtman [shehtmanrieger], see [Chagrov1997-cr, 8.8] or [bezhanishviliphdthesis, 3.2] for modern presentations ), which clarified the size and structure of irreducible elements of the free $n$-generated algebra, as well as some of its properties like completeness;
@@
Despite the importance of such advances, the situation presently in the literature presents a big contrast between what is known about Boolean algebras, distributive lattices and Heyting algebras: the former have descriptions of their free algebras, coproducts, and free one-step extensions for any number of generators, and any original algebra, whilst the latter are only available for special cases, namely those involving finite algebras.
The main contribution of this paper is to provide a generalization of the constructions presented by Ghilardi to the infinite case. The main hurdle in this is realizing wherein finiteness plays a role, to allow using Priestley duality in place of Birkhoff duality for finite distributive lattices, a task which turns out to not be obvious. This is equivalent to describing, in dual terms, the left adjoint to the inclusion of $\mathbf{HA}$, the category of Heyting algebras, in $\mathbf{DL}$, the category of distributive lattices. Due to the duality theory for such categories, this yields a right adjoint to the inclusion of $\mathbf{Esa}$, the category of Esakia spaces and continuous p-morphisms, into the category $\mathbf{Pries}$ of Priestley spaces and continuous order-preserving maps. Having this at hand, we employ it for several purposes:
@@
    We provide a description of the free Heyting algebras on any number of generators, and provide a concrete description of coproducts of Heyting algebras as well as of the pushout of Heyting algebras. From this we derive some results concerning co-distributivity of Heyting algebras.
@@
[…]
@@
    We first show that this construction can be adapted when dealing extensions of $\mathsf{IPC}$, or algebraically, subvarieties of $\mathbf{HA}$. Rather than providing a general theory, we focus on three important and illustrative cases: Boolean algebras, $\mathbf{KC}$ algebras and Gödel algebras. As a consequence of our analysis we show that $\mathsf{LC}$ has uniform local tabularity (first studied in [Shehtman2016]under the name finite formula depth property), and that in fact every formula is equivalent to one of implication depth $2$.
  4. 4.
    Finally, we look at the relationship between the category of posets with p-morphisms and the category of posets with monotone maps, and provide a right adjoint to the inclusion which is heavily inspired by the above constructions. The latter has found applications in the coalgebraic study of intuitionistic modal logic [coalgebraicintuitionisticmodal], addressing the open problem of how to represent Kripke frames for intuitionistic modal logics coalgebraically. We also show that the two constructions can be connected by the concept of an order-compactification.
The structure of the paper is as follows: in Section 2 we recall some basic preliminaries, and fix some notation moving forward. Section 3 contains the main technical tools, in the form of the description of the free Heyting algebra generated by a distributive lattice preserving a given set of relative pseudocomplements; in Section 4 we showcase some basic applications of the aforementioned theory to the study of Heyting algebras: some well-known results, like the amalgamation theorem, are shown here through direct methods, as well as some seemingly new results, like the comonadicity of the category of Heyting algebras over the category of distributive lattices. In Section 5 we look at an adjunction holding between Boolean algebras and Heyting algebras, related to the regularization functor, and show that the left adjoint of such a functor admits a description similar to the one given by the dual construction just given, and from this some applications to the study of inquisitive logic. In Section 6 we look at the case of subvarieties of Heyting algebras and illustrate how these constructions can be fruitfully analysed there. In Section 7 we look at the corresponding relationship between the category of posets with p-morphisms and the category of posets with monotone maps. We conclude in Section 8 with some outlines of future work.
@@ 2 Preliminaries
Proposition 2.7.
If $X$ is an arbitrary set, then $\mathbf{D}(X)$ is isomorphic to the free distributive lattice on $X$ many generators. Moreover, if $X$ is finite, then $F_{D}(X)$ is isomorphic to $(\mathcal{P}(X),\supseteq)$with the discrete topology.
@@
Proof.
Note that if each $X_{n}$ is an Esakia space, and each $p_{i,j}$ is a p-morphism, then the inverse limit is dual to the directed limit of distributive lattices; in these conditions, one can define a Heyting algebra structure on the directed union by taking the equivalence classes of the relative pseudocomplements arising in the finitely many stages, which will be well-defined since $p_{i,j}^{-1}$ are Heyting algebra homomorphisms. ∎
We also recall the well-known construction of the Vietoris hyperspace, due to Leopold Vietoris [Vietoris1921] (see also [guramremarksonhyperspacesforpriestleyspaces]). This can equivalently be given as follows: if $(X,\leq)$ is a Priestley space, let $V(X)$ be the set of all non-empty closed subsets of $X$. We give this set a topology by considering a subbasis consisting of, for $U,V\subseteq X$ clopen sets:
@@
This is sometimes called the “hit-and-miss topology”, and the resulting space is called the Vietoris hyperspace [^1] Vietoris’ original definition considers instead $[U]$ and $\langle V\rangle$ for $U$ open. But note that these topologies are equivalent when $X$ is Stone, since closed subsets are compact in such spaces.. On this space we define an order relation $C\preceq D$ if and ony if [^2] The reason this order is reverse inclusion has to do with the choice to use upsets rather than downsets; in [ghilardifreeheyting], the author uses downsets, which is why the swap no reversion is needed. $D\subseteq C$. Then we have the following:
Lemma 2.9.
The space $(V(X),\preceq)$ is a Priestley space. Indeed, it is an Esakia space.
@@
Otherwise, we have that $C\neq\emptyset$, so $C\in V(X)$, and since $C\notin\langle V_{j}\rangle$ for each $i$, by definition $C\in[U_{i}]$ for some $i$. Since $X-\bigcup_{i\in I}V_{i}\subseteq U_{i}$ then $X-U_{i}\subseteq\bigcup_{i\in I}V_{i}$, so by compactness, $X-U_{i}\subseteq V_{i_{0}}\cup...\cup V_{i_{n}}$. Hence consider the finite subcover $\{[U_{i}]\}\cup\{\langle V_{i_{j}}\rangle:j\in\{1,...,n\}\}$. If $D$ is any closed set, and $D\subseteq U_{i}$, then we are done; otherwise, $D\cap X-U_{i}\neq\emptyset$, so $D\cap V_{i_{0}}\cup...\cup V_{i_{n}}\neq\emptyset$, and so the result follows.
Now we prove the Priestley Separation axiom. Assume that $C\npreceq D$ are closed subsets. Since the space $X$ is a Stone space, we know that $C=\bigcap_{i\in I}V_{i}$ and $D=\bigcap_{j\in J}W_{j}$ where $V_{i},W_{j}$ are clopen. Hence there must be some $i$ such that $C\subseteq V_{i}$ and $D\nsubseteq V_{i}$; i.e., $C\in[V_{i}]$ and $D\notin[V_{i}]$. Since $[V_{i}]$ is clearly a clopen upset under this order, we verify the Priestley separation axiom.
Finally, consider an arbitrary clopen difference:
$Z=[U]\cap\langle V_{0}\rangle\cap...\cap\langle V_{n}\rangle.$
Then we want that ${\downarrow}Z$ is clopen. Indeed consider:
$\bigcap_{i=1}^{n}\langle U\cap V_{i}\rangle.$
If $C$ belongs to this set, then it contains a subset $D=\{x_{1},...,x_{n}\}$, elements in $V_{i}$ and $U$; so $D\subseteq U$, it is a closed subset, and intersects each of the sets. So $C\in{\downarrow}Z$. Conversely, if $C\in{\downarrow}Z$, then it is clear it belongs above. So since $\langle U\cap V_{i}\rangle$ are all subbasic clopens, we get that ${\downarrow}S$ is a clopen, as desired. ∎
@@ 3.1 Conceptual Idea of the Construction and the Finite Case
    $(a\vee b)\rightarrow c=(a\rightarrow c)\wedge(b\rightarrow c)$;
@@
    $a\rightarrow(b\wedge c)=(a\rightarrow b)\wedge(a\rightarrow c)$;
@@
    $(a\rightarrow b)\wedge(b\rightarrow c)\leq a\rightarrow c$.
@@ 3.2 G-Open Subsets and Vietoris Functors
Now assume that $p=f^{-1}$ is $g$-open, and $q=g^{-1}$. Assume that $f(x)\leq y$ where $x\in X$ and $y\in Y$. By duality, and abusing notation, this means that $p^{-1}[x]\subseteq y$. Consider $\mathsf{Fil}(x,\{p(q(a)):q(a)\in y\})$ and $\mathsf{Id}(\{p(q(b)):q(b)\notin y\})$. We claim these do not intersect. Because otherwise, for some $c\in x$, $c\wedge p(q(a))\leq p(q(b))$. Hence $c\leq p(q(a))\rightarrow p(q(b))$, since these exist and are preserved by $p$, and so $c\leq p(q(a)\rightarrow q(b))$. So $p(q(a)\rightarrow q(b))\in x$. Hence $q(a)\rightarrow q(b)\in p^{-1}[x]$, and so $q(a)\rightarrow q(b)\in y$, a contradiction. Hence by the Prime filter theorem (see Theorem 2.4), we can extend $x$ to a filter $x^{\prime}$ which does not intersect the presented ideal. By definition, working up to natural isomorphism, we then have that $g(f(x^{\prime}))=g(y)$, which was to show. ∎
@@
Throughout, fix $g:X\to Y$ a continuous and order-preserving map (the case for several such maps being preserved is entirely similar). Recall from Lemma 2.9 that $(V(X),\preceq)$ is a Priestley space. From this space we can move closer to the space we will be interested, by first considering $V_{r}(X)\subseteq V(X)$, the space of rooted closed subsets, with the induced order and the subspace topology. On this we can prove the following [^3] This is stated without proof in [Bezhanishvili2011heytingcoalgebra, Lemma 6.1]. The key idea of the proof below was communicated to me by Mamuka Jibladze.:
Lemma 3.4.
@@
Finally, we will refine the rooted subsets to the ones we are truly interested in – the $g$-open ones. Let $V_{g}(X)$ be the set of $g$-open, closd and rooted subsets. The following lemma encapsulates the key fact about this.
Lemma 3.5.
@@
Assume that $M$ is not $g$-open. This means that if $x$ is the root of $M$, there is some $y$ such that $x\leq y$, but for each $k\in M$, $g(k)\neq g(y)$. Hence $g(y)\notin g[M]$, and hence $\{g(y)\}\cap g[M]=\emptyset$. Since $g$ is a continuous map between Stone spaces, $g[M]$ is closed; hence, since this is a Priestley space, there is a clopen set of the form $-U_{i}\cup V_{i}$ for $U_{i},V_{i}$ clopen upsets, such that $g[M]\subseteq-U_{i}\cup V_{i}$, and $g(y)\notin-U_{i}\cup V_{i}$: indeed, one can find a clopen set $W$, and in these spaces we have that each clopen set can be written as $W=\bigcap_{i=1}^{n}(-U_{i}\cup V_{i})$. Then we have that $M\subseteq-g^{-1}[U_{i}]\cup g^{-1}[V_{i}]$ and $M\cap{\downarrow}(g^{-1}[U_{i}]-g^{-1}[V_{i}])\neq\emptyset$. Hence consider
@@
Because $X$ has relative pseudocomplements of sets of the form $g^{-1}[U]$, by Proposition 2.5 we have that ${\downarrow}(g^{-1}[U_{i}]-g^{-1}[V_{i}])$ will be clopen in $X$. Hence the set $S$ is an intersection of subbasic sets. Moreover, $M$ belongs there. Now if $N$ belongs there, then first $N\subseteq-g^{-1}[U_{i}]\cup g^{-1}[V_{i}]$; and also, by downwards closure, its root $k$ will be below some point $y^{\prime}$ in $g^{-1}[U_{i}]-g^{-1}[V_{i}]$. But then for each $m\in N$, $g(m)\neq g(y^{\prime})$, so $N$ is not $g$-open.